OpenStudy (anonymous):

What is the slope of the line represented by the equation 6x –3y = 4 ?

OpenStudy (anonymous):

To get the slope change it into slope intercept form 6x - 3y = 4 -6x -6x -3y = -6x + 4 -3y/-3 = -6/-3 + 4/-3 y = 2x - 1.3 So can you tell me what the slope is?

OpenStudy (anonymous):

y = mx + b When m = slope and b = y-intercept

OpenStudy (anonymous):

@Ashbtr18 are you there?